### v2

@@ 3.3 MaxEnt RL with Diffusion Policy
Action Selection for inference. Previous studies [36, 19, 37, 38] have found that a deterministic policy typically outperforms its stochastic counterpart during testing. Consequently, we employ an action selection technique to further refine the policy after training. Specifically, $M$ action candidates are sampled from the diffusion policy, and the action with the highest Q-value is selected to interact with the RL environment.
@@ 4 Related Work
MaxEnt RL A variety of approaches have been proposed to achieve the MaxEnt RL objective. SQL [9] introduces soft Q-learning to learn the optimal soft Q-function and trains an energy-based model using the amortized Stein variational gradient method to generate actions according to the exponential of the optimal soft Q-function. SAC [10] presents the soft actor-critic algorithm, which iteratively improves the policy towards a higher soft value, and provides an implementation using Gaussian policies. MEow [36] employs energy-based normalizing flows as unified policies to represent both the actor and the critic, simplifying the training process for MaxEnt RL. This paper highlights the importance of policy representation within the MaxEnt RL framework: a more expressive policy representation enhances exploration and facilitates closer convergence to the optimal MaxEnt policy. Diffusion models, which are more expressive than Gaussian distributions and energy-based normalizing flows and easier to train and sample than energy-based models, present an ideal policy representation that effectively balances expressiveness and the complexity of training and inference.
@@
Diffusion Policies for Offline RL. Offline RL attempts to learn a well-performing policy from a pre-collected dataset. Collected by multiple policies, the offline datasets may exhibit high skewness and multi-modality. Diffusion Policy [18] trains a diffusion model to approximate the multi-modal expert behavior by behavior cloning. To optimize the policy for higher performance, Diffusion-QL [19] combines the diffusion loss with Q-value loss evaluated on the generated actions, CEP [39] trains a separate guidance network using Q-function to guide the actions to regions with high Q values, and EDA [38] employs direct preference optimization to align the diffusion policy with Q-function. To improve the training and inference speed of diffusion policy, EDP [40] adopts action approximation and efficient ODE sampler DPM-solver for action generation, and CPQL [41] utilizes the consistency policy [42], a one-step diffusion policy. Due to the lack of online samples, the above approaches require staying close to the behavior policy to prevent out-of-distribution actions whose performances are unpredictable. However, in this paper, we focus on online RL, where online interactions are accessible to correct the errors in value evaluation. Therefore, different techniques should be developed to employ diffusion models in online RL.
@@
Diffusion Policies for Online RL. In online RL, a key challenge lies in balancing exploration and exploitation. Previous studies [32, 43, 44, 45] apply expressive diffusion models as policy representations to promote the exploration of the state-action space. QSM [32] fits the exponential of the Q-function by training a score network to approximate the action gradient of the Q-function. DIPO [43] improves the actions by applying the action gradient of the Q-function and clones the improved actions. QVPO [44] weights the diffusion loss with the Q-value, assigning probabilities to actions that are linearly proportional to the Q-value. DACER [45] optimizes the Q-value loss to generate actions with high Q values and adds extra noise to the generated actions to keep a constant policy entropy. Unlike previous approaches, we employ the MaxEnt RL objective to encourage exploration and enhance policy robustness. Similar to QSM, we train the diffusion model to fit the exponential of the Q-function. However, our Q-weighted noise estimation method is more accurate and stable. Furthermore, we include policy entropy when computing the Q-function, which can further promote exploration.
@@ 5.1 A Toy Multi-goal Environment
In this subsection, we use a 2-D multi-goal environment [9] to demonstrate the effectiveness of MaxEntDP. In this environment, the agent is a 2-D point mass trying to reach one of four symmetrically placed goals. The state and action are position and velocity, respectively. And the reward is a penalty for the velocity and distance from the closest goal. Under the MaxEnt RL objective, the optimal policy is to choose one goal randomly and then move toward it. Figure 1 shows the trajectories generated by the diffusion policy during the training process. We can see that MaxEntDP can effectively explore the state-action space and learn a multi-modal policy that approaches the optimal MaxEnt policy.
@@ 5.2 Comparative Evaluation
Policy Representations. To reveal the superiority of applying diffusion models as policy representations to achieve the MaxEnt RL objective, we compare the performance of MaxEntDP on Mujoco benchmarks [46] with other algorithms. Our chosen baseline algorithms include SAC [36], MEow [36], and TD3 [47]. SAC and MEow are two methods to pursue the same MaxRnt RL objective using Gaussian policy and energy-based normalizing flow policy, and TD3 provides a contrast to the deterministic policy. Figure 2 shows that MaxEntDP surpasses (a-d) or matches (e-f) baseline algorithms on all tasks, and its evaluation variance is much smaller than other algorithms. This result indicates that the combination of MaxEntRL and diffusion policy effectively balances exploration and exploitation, enabling rapid convergence to a robust and high-performing policy.
Diffusion Model Training Methods. In this subsection, we demonstrate the advantages of the proposed Q-weighted Noise Estimation method (QNE) on training diffusion models, compared to two competing methods, QSM and iDEM. We replace the QNE module with QSM and iDEM to observe performance differences. As shown in Figure 3(a), the performance of QSM and iDEM improves initially but then fluctuates after reaching a high level. This may be due to both methods relying on the gradient computation of the Q-function to estimate the score function. When the Q-value is large, its gradient typically varies much across different actions, leading to a high variance in score function estimation for QSM and iDEM, as illustrated in Figure 3(b). This increased variance causes instability in the training of the noise prediction network. In contrast, QNE exhibits significantly lower variance, and its performance improves steadily throughout the training process.
@@ 5.3 Ablation analysis
In addition, we analyze the function of the MaxEnt RL objective by removing the probability approximation module in MaxEntDP. After doing this, we compute the original Q-function rather than the soft Q-function in the policy evaluation step. As shown in Figure 5, the performance decreases and exhibits greater variance after excluding policy entropy in the Q-function. This implies that the MaxEnt RL objective can benefit policy learning: it not only encourages the action distribution at the current step to be more stochastic (by fitting the exponential of Q-function), but also encourages transferring to the states with higher policy entropy (by computing the soft Q-function). Therefore, the MaxEnt RL objective shows a stronger exploration ability of the whole state-action space, leading to an efficient and stable training process.
@@ A.1 Proofs for Soft Actor-Critic algorithm
which proves that the soft Bellman operator $\mathcal{T}^{\pi}$ is a contraction. It has a unique fixed point $Q^{\pi}$. When Q-function loss $L(\theta)=0$, the Q-function $Q_{\theta}$ satisfies the soft Bellman equation $Q_{\theta}(\boldsymbol{s},\boldsymbol{a})=r(\boldsymbol{s},\boldsymbol{a})+\gamma\mathbb{E}_{\boldsymbol{s}^{\prime}\sim p,\boldsymbol{a}^{\prime}\sim\pi_{\phi}}\left[Q_{\theta}(\boldsymbol{s}^{\prime},\boldsymbol{a}^{\prime})-\beta\log\pi_{\phi}(\boldsymbol{a}^{\prime}|\boldsymbol{s}^{\prime})\right]$ for all $(\boldsymbol{s},\boldsymbol{a})\in|\mathcal{S}|\times|\mathcal{A}|$, indicating that $Q_{\theta}$ converges to the true soft Q-function of $\pi_{\phi}$. ∎
Lemma 8.
@@
Since the new policy $\pi_{\phi_{k+1}}$ is the minimizer of the loss defined in Equation 3, it holds that
@@
$\hskip 46.25275pt+\gamma^{2}\mathbb{E}_{\boldsymbol{s}_{t+1}\sim p,\boldsymbol{a}_{t+1}\sim\pi_{\phi_{k+1}},\boldsymbol{s}_{t+2}\sim p,\boldsymbol{a}_{t+2}\sim\pi_{\phi_{k}}}\left[Q^{\pi_{\phi_{k}}}(\boldsymbol{s}_{t+2},\boldsymbol{a}_{t+2})-\beta\log\pi_{\phi_{k}}(\boldsymbol{a}_{t+2}|\boldsymbol{s}_{t+2})\right]$ (41)
@@
which is proved by repeatedly expanding $Q^{\pi_{\phi_{k}}}$ using the soft Bellman equation and applying Equation 37. Then the proof for Lemma 8 is completed. ∎
Theorem 9.
@@
According to Lemma 7 and 8, when $L(\theta_{k})=0$ and $L(\phi_{k})$ be minimized for each $k$, we have $\forall k,Q^{\pi_{\phi_{k+1}}}\geq Q^{\pi_{\phi_{k}}}$. This indicates that the sequence $Q^{\pi_{\phi_{k}}}$ is monotonically increasing. Furthermore, the Q-function is bounded since both the reward and entropy are bound. Therefore, when $k\to\infty$, the policy sequence converges to some $\pi^{*}$. Below we will prove that $\pi^{*}$ is the optimal MaxEnt policy within $\pi$.
@@ A.4 The derivation of the iDEM method
We provide the derivation of the iDEM method to demonstrate the difference and relationship between Q-weighted noise estimation and iDEM. Our derivation is equivalent to the official proof of iDEM although in a different way.
@@
$\nabla_{\boldsymbol{a}_{t}}\log p(\boldsymbol{a}_{t})$ $=\frac{1}{\sqrt{\sigma(\alpha_{t})}}\int\nabla_{\boldsymbol{a}_{0}}w(\boldsymbol{a}_{0})\cdot\mathcal{N}(\boldsymbol{a}_{0}|\frac{1}{\sqrt{\sigma(\alpha_{t})}}\boldsymbol{a}_{t},\frac{\sigma(-\alpha_{t})}{\sigma(\alpha_{t})}\boldsymbol{I})\text{d}\boldsymbol{a}_{0}.$ (59)
@@ A.5 Probabity Approximation of Diffusion Policy using Numerical Integration Techniques
where $c=-\frac{d}{2}\log(2\pi e)+\frac{d}{2}\int_{-\infty}^{+\infty}\sigma(\alpha_{t})\text{d}\alpha_{t}$ with $d$ being the dimension of $\boldsymbol{a}_{0}$, $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$, $\boldsymbol{a}_{t}=\sqrt{\sigma(\alpha_{t})}\boldsymbol{a}_{0}+\sqrt{\sigma(-\alpha_{t})}\boldsymbol{\epsilon}$. First, we replace the integral variable for $\sigma(\alpha_{t})$ as it has a narrow integration domain of $[0,1]$:
@@
Appendix B Experimental Details.
B.1 Hyperparameter Analysis

### v3

@@ 3.3 MaxEnt RL with Diffusion Policy
Action Selection for Inference. Previous studies (Chao et al., 2024; Wang et al., 2023; Mao et al., 2024; Chen et al., 2024b) have found that a deterministic policy typically outperforms its stochastic counterpart during testing. Consequently, we employ an action selection technique to further refine the policy after training. Specifically, $M$ action candidates are sampled from the diffusion policy, and the action with the highest Q-value is selected to interact with the RL environment.
@@ 4 Related Work
MaxEnt RL A variety of approaches have been proposed to achieve the MaxEnt RL objective. SQL (Haarnoja et al., 2017) introduces soft Q-learning to learn the optimal soft Q-function and trains an energy-based model using the amortized Stein variational gradient method to generate actions according to the exponential of the optimal soft Q-function. SAC (Haarnoja et al., 2018) presents the soft actor-critic algorithm, which iteratively improves the policy towards a higher soft value, and provides an implementation using Gaussian policies. To improve the sample efficiency of SAC, CrossQ (Bhatt et al., 2024) and BRO (Nauman et al., 2024) construct larger critic networks and apply a suite of regularization techniques to stabilize training. MEow (Chao et al., 2024) employs energy-based normalizing flows as unified policies to represent both the actor and the critic, simplifying the training process for MaxEnt RL. This paper highlights the importance of policy representation within the MaxEnt RL framework: a more expressive policy representation enhances exploration and facilitates closer convergence to the optimal MaxEnt policy. Diffusion models, which are more expressive than Gaussian distributions and energy-based normalizing flows and easier to train and sample than energy-based models, present an ideal policy representation that effectively balances expressiveness and the complexity of training and inference.
@@
Diffusion Policies for Offline RL. Offline RL attempts to learn a well-performing policy from a pre-collected dataset. Collected by multiple policies, the offline datasets may exhibit high skewness and multi-modality. Diffusion Policy (Chi et al., 2023) trains a diffusion model to approximate the multi-modal expert behavior by behavior cloning. To optimize the policy for higher performance, Diffusion-QL (Wang et al., 2023) combines the diffusion loss with Q-value loss evaluated on the generated actions, CEP (Lu et al., 2023) trains a separate guidance network using Q-function to guide the actions to regions with high Q values, and EDA (Chen et al., 2024b) employs direct preference optimization to align the diffusion policy with Q-function. To improve the training and inference speed of diffusion policy, EDP (Kang et al., 2024) adopts action approximation and efficient ODE sampler DPM-solver for action generation, and CPQL (Chen et al., 2024c) utilizes the consistency policy (Song et al., 2023), a one-step diffusion policy. Due to the lack of online samples, the above approaches require staying close to the behavior policy to prevent out-of-distribution actions whose performances are unpredictable. However, in this paper, we focus on online RL, where online interactions are accessible to correct the errors in value evaluation. Therefore, different techniques should be developed to employ diffusion models in online RL.
@@
Diffusion Policies for Online RL. In online RL, a key challenge lies in balancing exploration and exploitation. Previous studies (Psenka et al., 2024; Yang et al., 2023; Ding et al., 2024; Wang et al., 2024) apply expressive diffusion models as policy representations to promote the exploration of the state-action space. QSM (Psenka et al., 2024) fits the exponential of the Q-function by training a score network to approximate the action gradient of the Q-function. DIPO (Yang et al., 2023) improves the actions by applying the action gradient of the Q-function and clones the improved actions. QVPO (Ding et al., 2024) weights the diffusion loss with the Q-value, assigning probabilities to actions that are linearly proportional to the Q-value. DACER (Wang et al., 2024) optimizes the Q-value loss to generate actions with high Q values and adds extra noise to the generated actions to keep a constant policy entropy. Unlike previous approaches, we employ the MaxEnt RL objective to encourage exploration and enhance policy robustness. Similar to QSM, we train the diffusion model to fit the exponential of the Q-function. However, our Q-weighted noise estimation method is more accurate and stable. Furthermore, we include policy entropy when computing the Q-function, which can further promote exploration.
@@ 5.1 A Toy Multi-goal Environment
In this subsection, we use a 2-D multi-goal environment (Haarnoja et al., 2017) to demonstrate the effectiveness of MaxEntDP. In this environment, the agent is a 2-D point mass trying to reach one of four symmetrically placed goals. The state and action are position and velocity, respectively. And the reward is a penalty for the velocity and distance from the closest goal. Under the MaxEnt RL objective, the optimal policy is to choose one goal randomly and then move toward it. Figure 2 shows the trajectories generated by the diffusion policy during the training process. We can see that MaxEntDP can effectively explore the state-action space and learn a multi-modal policy that approaches the optimal MaxEnt policy.
@@ 5.2 Comparative Evaluation
Policy Representations. To reveal the superiority of applying diffusion models as policy representations to achieve the MaxEnt RL objective, we compare the performance of MaxEntDP on Mujoco benchmarks (Todorov et al., 2012) with other algorithms. Our chosen baseline algorithms include SAC (Haarnoja et al., 2018), MEow (Chao et al., 2024), and TD3 (Fujimoto et al., 2018). SAC and MEow are two methods to pursue the same MaxRnt RL objective using Gaussian policy and energy-based normalizing flow policy, and TD3 provides a contrast to the deterministic policy. Figure 1 shows that MaxEntDP surpasses (a-d) or matches (e-f) baseline algorithms on all tasks, and its evaluation variance is much smaller than other algorithms. This result indicates that the combination of MaxEntRL and diffusion policy effectively balances exploration and exploitation, enabling rapid convergence to a robust and high-performing policy.
Diffusion Models Training Methods. In this subsection, we demonstrate the advantages of the proposed Q-weighted Noise Estimation method (QNE) on training diffusion models, compared to two competing methods, QSM and iDEM. We replace the QNE module with QSM and iDEM to observe performance differences. As shown in Figure 3(a), the performance of QSM and iDEM improves initially but then fluctuates after reaching a high level. This may be due to both methods relying on the gradient computation of the Q-function to estimate the score function. When the Q-value is large, its gradient typically varies much across different actions, leading to a high variance in score function estimation for QSM and iDEM, as illustrated in Figure 3(b). This increased variance causes instability in the training of the noise prediction network. In contrast, QNE exhibits significantly lower variance, and its performance improves steadily throughout the training process.
@@ 5.3 Ablation Analysis
In addition, we analyze the function of the MaxEnt RL objective by removing the probability approximation module in MaxEntDP. After doing this, we compute the original Q-function rather than the soft Q-function in the policy evaluation step. As shown in Figure 5, the performance decreases and exhibits greater variance after excluding policy entropy in the Q-function. This implies that the MaxEnt RL objective can benefit policy learning: it not only encourages the action distribution at the current step to be more stochastic (by fitting the exponential of Q-function), but also encourages transferring to the states with higher entropy (by computing the soft Q-function). Therefore, the MaxEnt RL objective shows a stronger exploration ability of the whole state-action space, leading to an efficient and stable training process.
@@ 6 Conclusion
Acknowledgements
We extend our gratitude to the ICML reviewers who evaluated MaxEntDP for their valuable insights and feedback. This work was supported by the National Key Research and Development Program of China (No. 2021ZD0201504).
@@ A.1 Proofs for Soft Actor-Critic Algorithm
which proves that the soft Bellman operator $\mathcal{T}^{\pi}$ is a contraction. It has a unique fixed point $Q^{\pi}$. When Q-function loss $L(\theta)=0$, the Q-function $Q_{\theta}$ satisfies the soft Bellman equation $Q_{\theta}(\boldsymbol{s},\boldsymbol{a})=r(\boldsymbol{s},\boldsymbol{a})+\gamma\mathbb{E}_{\boldsymbol{s}^{\prime}\sim p,\boldsymbol{a}^{\prime}\sim\pi_{\phi}}\left[Q_{\theta}(\boldsymbol{s}^{\prime},\boldsymbol{a}^{\prime})-\beta\log\pi_{\phi}(\boldsymbol{a}^{\prime}|\boldsymbol{s}^{\prime})\right]$ for all $(\boldsymbol{s},\boldsymbol{a})\in|\mathcal{S}|\times|\mathcal{A}|$, indicating that $Q_{\theta}$ converges to the true soft function of $\pi_{\phi}$. ∎
Lemma A.2.
@@
Since the new policy $\pi_{\pi_{k+1}}$ is the minimizer of the loss defined in Equation 3, it holds that
@@
$\hskip 49.14307pt+\gamma^{2}\mathbb{E}_{\boldsymbol{s}_{t+1}\sim p,\boldsymbol{a}_{t+1}\sim\pi_{\phi_{k+1}},\boldsymbol{s}_{t+2}\sim p,\boldsymbol{a}_{t+2}\sim\pi_{\phi_{k}}}\left[Q^{\pi_{\phi_{k}}}(\boldsymbol{s}_{t+2},\boldsymbol{a}_{t+2})-\beta\log\pi_{\phi_{k}}(\boldsymbol{a}_{t+2}|\boldsymbol{s}_{t+2})\right]$ (41)
@@
which is proved by repeatedly expanding $Q^{\pi_{\phi_{k}}}$ using the soft Bellman equation and applying Equation 37. Then the proof for Lemma A.2 is completed. ∎
Theorem A.3.
@@
According to Lemma A.1 and A.2, when $L(\theta_{k})=0$ and $L(\phi_{k})$ be minimized for each $k$, we have $\forall k,Q^{\pi_{\phi_{k+1}}}\geq Q^{\pi_{\phi_{k}}}$. This indicates that the sequence $Q^{\pi_{\phi_{k}}}$ is monotonically increasing. Furthermore, the Q-function is bounded since both the reward and entropy are bound. Therefore, when $k\to\infty$, the policy sequence converges to some $\pi^{*}$. Below we will prove that $\pi^{*}$ is the optimal MaxEnt policy within $\pi$.
@@ A.4 The Derivation of the iDEM Method
We provide the derivation of the iDEM method to demonstrate the difference and relationship between Q-weighted noise estimation and iDEM. Our derivation is equivalent to the official proof of iDEM, although in a different way.
@@
$\nabla_{\boldsymbol{a}_{t}}\log p(\boldsymbol{a}_{t})$ $=\frac{1}{\sqrt{\sigma(\alpha_{t})}}\int\left(\nabla_{\boldsymbol{a}_{0}}w(\boldsymbol{a}_{0})\right)\cdot\mathcal{N}(\boldsymbol{a}_{0}|\frac{1}{\sqrt{\sigma(\alpha_{t})}}\boldsymbol{a}_{t},\frac{\sigma(-\alpha_{t})}{\sigma(\alpha_{t})}\boldsymbol{I})\text{d}\boldsymbol{a}_{0}.$ (59)
@@ A.5 Probabity Approximation of Diffusion Policy Using Numerical Integration Techniques
where $c=-\frac{d}{2}\log(2\pi e)+\frac{d}{2}\int_{-\infty}^{+\infty}\sigma(\alpha_{t})\text{d}\alpha_{t}$ with $d$ being the dimension of $\boldsymbol{a}_{0}$, $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$, $\boldsymbol{a}_{t}=\sqrt{\sigma(\alpha_{t})}\boldsymbol{a}_{0}+\sqrt{\sigma(-\alpha_{t})}\boldsymbol{\epsilon}$. First, using the equation $\alpha_{t}=\log\frac{\sigma(\alpha_{t})}{1-\sigma(\alpha_{t})}$, we change the integral variable from $\alpha_{t}$ to $\sigma(\alpha_{t})$ as $\sigma(\alpha_{t})$ has a narrow integration domain of $(0,1)$:
@@
Appendix B Supplementary Related Work on Diffusion-based Energy Models
A line of work focuses on applying the expressive diffusion models to approximate the exponential of a given energy function. QSM (Psenka et al., 2024) trains the score function by aligning it with the gradient of the energy function. iDEM (Akhound-Sadegh et al., 2024) proposes a weighted sum of the gradient of the energy function to estimate the true score function. These two approaches, which are based on gradient computation, suffer from a large estimation variance and demonstrate training instability when used for diffusion policy optimization, as evidenced in Section 5.2. Recently, a work (Sendera et al., 2024) considers the Euler-Maruyama samplers of diffusion models as continuous generative flow networks (GFlowNets), and exploits the trajectory balance objective to train diffusion models. In this method, a replay buffer is used to store the sample generation trajectories of diffusion models, which may cause a heavy memory burden. The model-based diffusion (Pan et al., 2024) proposes the Monte Carlo estimation for computing the score function and uses the Monte Carlo score ascent to generate samples following the Boltzmann distribution of a given function. The model-based diffusion is similar to our QNE method, however, QNE has several properties that matter in RL training:
  • We use a parameterized network to approximate the scaled score function, while the model-based diffusion needs to compute the score function using Monte Carlo estimation when generating samples. Therefore, sample generation of model-based diffusion is time-consuming, which will slow the training speed of the RL algorithms.
  • We adopt ancestral sampling in DDPM to generate samples, that are more diverse than that of Monte Carlo score ascent used in model-based diffusion.
  • We propose to modify the standard Gaussian to the truncated Gaussian in QNE to model the action distribution with a bounded action space. However, model-based diffusion can not address such a bounded distribution.
These properties make QNE well-suited for the optimization of diffusion policy.
Appendix C Experimental Details
C.1 Hyperparameters Settings
All experiments in this paper are conducted on a GPU of Nvidia GeForce RTX 3090 and a CPU of AMD EPYC 7742. Our implementation of SAC, MEow, TD3, QSM, DACER, QVPO, and DIPO follows their official codes: https://github.com/toshikwa/soft-actor-critic.pytorch, https://github.com/ChienFeng-hub/meow, https://github.com/sfujim/TD3, https://github.com/Alescontrela/score_matching_rl, https://github.com/happy-yan/DACER-Diffusion-with-Online-RL, https://github.com/wadx2019/qvpo, and https://github.com/BellmanTimeHut/DIPO. The shared hyperparameters of all algorithms are listed in Table 1222When comparing with other diffusion-based algorithms, MaxEntDP uses 3-layer MLPs as the actor and critic networks following the default settings of these algorithms. In other experiments, 2-layer MLPs are used as they can already attain good performance..
Table 1: The shared hyperparameters of all algorithms.
Hyperparameter MaxEntDP SAC MEow TD3 QSM DACER QVPO DIPO
Batch size 256 256 256 256 256 256 256 256
Discount factor $\gamma$ 0.99 0.99 0.99 0.99 0.99 0.99 0.99 0.99
No. of hidden layers 2 2 2 2 3 3 3 3
No. of hidden nodes 256 256 256 256 256 256 256 256
Actor learning rate 3e-4 3e-4 3e-4 3e-4 3e-4 3e-4 3e-4 3e-4
Critic learning rate 3e-4 3e-4 3e-4 3e-4 3e-4 3e-4 3e-4 3e-4
Activation mish relu relu relu mish mish mish mish
Replay buffer size 1e6 1e6 1e6 1e6 1e6 1e6 1e6 1e6
Diffusion steps 20 N/A N/A N/A 20 20 20 20
Action candidate number 10 N/A N/A N/A N/A N/A 32 N/A
C.2 Training Time
The training time for all algorithms is presented in Table 2. Leveraging the computation efficiency of JAX (Frostig et al., 2018) and the parallel processing capabilities of GPU, MaxEntDP demonstrates high training efficiency compared to competing methods, only behind TD3 and QSM. This highlights its advantage for real-world applications that require high computation efficiency.
Algorithm (2-layer MLP) MaxEntDP (jax) SAC MEow TD3
Training time (h) 3.9 4.6 11 1.7
Training time (h) 5.5 1.9 5.9 22.6 55.6
Appendix D Supplementary Experiments
D.1 Hyperparameter Analysis